### v1

# Approximation by continuous functions and its applications

Abstract
We prove that for every normal topological space $X$ and any function      $f:X\to\mathbb{R}$ there is a continuous function $g:X\to\mathbb{R}$ such that
$|f(x)-g(x)|\leq\frac{1}{2}\sup\limits_{p\in X}\inf\limits_{O(p)}\sup\limits_{a,b\in O(p)}|f(a)-f(b)|$
for all $x\in X$.
As an application of this result we prove the following statements to types of tightness in a space $Q_{p}(X,\mathbb{R})$ of all quasicontinuous real-valued functions with the topology $\tau_{p}$ of pointwise convergence: the countability of tightness (fan-tightness, strong fan-tightness) at a point $f$ of space $Q_{p}(X,\mathbb{R})$ implies the countability of tightness (fan-tightness, strong fan-tightness) of space $Q_{p}(X,Y)$ of all quasicontinuous functions from $X$ into any non-one-point metrizable space $Y$. This result is the answer to the open question in the class of metrizable spaces.
keywords:
approximation, quasicontinuous function, Lusin space, open Whyburn space, tightness, fan-tightness, strong fan-tightness, selection principle, continuous selection
MSC:
[2020] 54C35, 54A25, 54D20, 54C10
journal:
1 Introduction
A function $f:X\rightarrow Y$ is quasicontinuous at $x$ if for any open set $V$ containing $f(x)$ and any $U$ open containing $x$, there exists a nonempty open set $W\subseteq U$ such that $f(W)\subseteq V$. It is quasicontinuous if it is quasicontinuous at every point. Call a set semi-open (or quasi-open) if it is contained in the closure of its interior. Then $f:X\rightarrow Y$ is quasicontinuous if and only if the inverse of every open set is quasi-open.
Let $X$ and $Y$ be Hausdorff topological spaces, $Q_{p}(X,Y)=(Q(X,Y),\tau_{p})$ be the space of all quasicontinuous functions on $X$ with values in $Y$ and $\tau_{p}$ be the pointwise convergence topology.
We were motivated for this paper by the problem of tightness in a space of quasi-continuous functions.
It is an open problem whether the space $Q_{p}(X,\mathbb{R})$ is topologically homogeneous (note that this space is not a linear space, since the sum of quasi-continuous functions is not necessarily quasi-continuous). In particular, it is unknown whether the points of $Q_{p}(X,\mathbb{R})$ can be distinguished using some topological property or cardinal invariant. In this regard, in (Osipov , Question 1) the second author raised a question on tightness in the space $Q_{p}(X,\mathbb{R})$. It was proved in Osipov that for every continuous $f:X\to\mathbb{R}$ we have $t(f,Q_{p}(X,\mathbb{R}))=\omega$ if and only if the space $X$ is $\mathcal{K}_{\Omega}$-Lindelof, and it was unknown whether the same is true for non-continuous functions.
In this paper we prove that the property of countable tightness (as well as countable fan-tightness and countable strong fan-tightness) can not distinguish points of the space $Q_{p}(X,Y)$ whenever $Y$ is metrizable.
One of the main tools of our proof is Theorem 3.7 below on approximation of arbitrary functions by continuous functions (in the metric of uniform convergence). Although its formulation looks quite classical, the authors were unable to find this statement in the literature. Section 3 is devoted to the proof of this theorem.
In Sections 4 and 5, we prove some other technical lemmas needed for Section 6, where we finally investigate the properties of countable tightness, fan-tightness and strong fan-tightness in $Q_{p}(X,Y)$.
2 Preliminaries
Let us recall some properties of a topological space $X$.
(1) A space $X$ has countable tightness at a point $x$ (denoted $t(x,X)=\omega$) if $x\in\overline{A}$, then $x\in\overline{B}$ for some countable $B\subseteq A$. A space $X$ has countable tightness (denoted $t(X)=\omega$) if $t(x,X)=\omega$ for every $x\in X$.
(2) A space $X$ has countable fan-tightness at a point $x$ (denoted $vet(x,X)=\omega$) if for any countable family $\{A_{n}:n\in\omega\}$ of subsets of $X$ satisfying $x\in\bigcap_{n\in\omega}\overline{A_{n}}$ it is possible to select finite sets $K_{n}\subset A_{n}$ in such a way that $x\in\overline{\bigcup_{n\in\omega}K_{n}}$. A space $X$ has countable fan-tightness (denoted $vet(X)=\omega$) if $vet(x,X)=\omega$ for every $x\in X$.
(3) A space $X$ is said to have countable strong fan-tightness at a point $x$ (denoted $vet_{1}(x,X)=\omega$) if for each countable family $\{A_{n}:n\in\omega\}$ of subsets of $X$ such that $x\in\bigcap_{n\in\omega}\overline{A_{n}}$, there exist $a_{i}\in A_{i}$ such that $x\in\overline{\{a_{i}:i\in\omega\}}$. A space $X$ has countable strong fan-tightness (denoted $vet_{1}(X)=\omega$) if $vet_{1}(x,X)=\omega$ for every $x\in X$.
(4) A space $X$ is said to be open Whyburn if for every open set $A\subset X$ and every $x\in\overline{A}\setminus A$ there is an open set $B\subseteq A$ such that $\overline{B}\setminus A=\{x\}$ Os1 .
Note that the class of open Whyburn spaces is quite wide; for example, it includes all first countable regular spaces Os1 and, therefore, all metrizable spaces.
Let $X$ be a Tychonoff topological space, $C(X,\mathbb{R})$ be the space of all continuous functions on $X$ with values in $\mathbb{R}$ and $\tau_{p}$ be the pointwise convergence topology. Denote by $C_{p}(X,\mathbb{R})$ the topological space $(C(X,\mathbb{R}),\tau_{p})$.
The symbol $\bf{0}$ stands for the constant function to 00. A basic open neighborhood of $\bf{0}$ in $\mathbb{R}^{X}$ is of the form $[F,(-\varepsilon,\varepsilon)]=\{f\in\mathbb{R}^{X}:f(F)\subset(-\varepsilon,\varepsilon)\}$, where $F\in[X]^{<\omega}$ and $\varepsilon>0$.
Let us recall that a cover $\mathcal{U}$ of a set $X$ is called
$\bullet$ an $\omega$-cover if each finite set $F\subseteq X$ is contained in some $U\in\mathcal{U}$;
$\bullet$ a $\gamma$-cover if for any $x\in X$ the set $\{U\in\mathcal{U}:x\not\in U\}$ is finite.
In this paper $\mathcal{A}$ and $\mathcal{B}$ will be collections of the following covers of a space $X$:
$\mathcal{O}^{s}$ : the collection of all semi-open covers of $X$.
$\Omega$ : the collection of open $\omega$-covers of $X$.
$\mathcal{K}$: the collection $\mathcal{U}$ of open subsets of $X$ such that $X=\bigcup\{\overline{U}:U\in\mathcal{U}\}$.
$\mathcal{K}_{\Omega}$ is the set of $\mathcal{U}$ in $\mathcal{K}$ such that no element of $\mathcal{U}$ is dense in $X$, and for each finite set $F\subseteq X$, there is a $U\in\mathcal{U}$ such that $F\subseteq\overline{U}$.
A space is $\mathcal{K}_{\Omega}$-Lindelöf if each element of $\mathcal{K}_{\Omega}$ has a countable subset in $\mathcal{K}_{\Omega}$ Sh2 .
Definition 2.1.
(Kun ) A Hausdorff space $X$ is called a Lusin space (in the sense of Kunen) if
(a) Every nowhere dense set in $X$ is countable;
(b) $X$ has at most countably many isolated points;
(c) $X$ is uncountable.
If $X$ is an uncountable Hausdorff space then $X$ is $\mathcal{O}^{s}$-Lindelöf (semi-Lindelöf) if and only if $X$ is a Lusin space (Corollary 2.5 in GJR ).
If $X$ is a Lusin space, $X$ is hereditarily Lindelöf (Lemma 1.2 in Kun ). Hence, if $X$ is a regular Lusin space then $X$ is perfect normal (3.8.A.(b) in Eng ).
If $X$ is a Lusin space, so is every uncountable subspace (Lemma 1.1 in Kun ).
Many topological properties are defined or characterized in terms of the following classical selection principles (see H1 ). Let $\mathcal{A}$ and $\mathcal{B}$ be sets consisting of families of subsets of an infinite set $X$. Then:
$S_{1}(\mathcal{A},\mathcal{B})$ is the selection hypothesis: for each sequence $\{A_{n}:n\in\mathbb{N}\}$ of elements of $\mathcal{A}$ there is a sequence $\{b_{n}\}_{n\in\mathbb{N}}$ such that for each $n$, $b_{n}\in A_{n}$, and $\{b_{n}:n\in\mathbb{N}\}\in\mathcal{B}$.
$S_{fin}(\mathcal{A},\mathcal{B})$ is the selection hypothesis: for each sequence $\{A_{n}:n\in\mathbb{N}\}$ of elements of $\mathcal{A}$ there is a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of finite sets such that for each $n$, $B_{n}\subseteq A_{n}$, and $\bigcup_{n\in\mathbb{N}}B_{n}\in\mathcal{B}$.
In this paper, by metric space we mean a non-trivial metric space, that is, of cardinality greater than 1.
For other notation and terminology almost without exceptions we follow the Engelking’s book Eng and the papers Os1 ; Osipov .
3 Approximation by continuous functions
Our goal in this section is to prove Theorem 3.7 and its corollary Theorem 3.9 below. The notion $\mathrm{dc}$ in its statement is defined as follows.
Definition 3.1.
Let $X$ be a topological space and let $(Y,d)$ be a metric space. For all functions $f:X\to Y$ we use the following notation:
  • 1.
    $\mathrm{dc}(f,E)=\sup\limits_{x,y\in E}d(f(x),f(y))$ for all $E\subseteq X$;
  • 2.
    $\mathrm{dc}(f,x)=\inf\limits_{O(x)}\mathrm{dc}(f,O(x))$;
  • 3.
    $\mathrm{dc}(f)=\sup\limits_{x\in X}\mathrm{dc}(f,x)$.
Our main tool in this section is Michael’s Selection Theorem. Before we formulate it, we must recall a few notions from HHM .
Definition 3.2.
Suppose that $X$ and $Y$ are topological spaces. A set-valued mapping $\Phi:X\to 2^{Y}$ is called lower semicontinuous, if for every $x_{0}\in X$, $y_{0}\in\Phi(x_{0})$ and a neighborhood $V$ of $y_{0}$ there is a neighborhood $U$ of $x_{0}$ such that for all $x\in U$ we have $\Phi(x)\cap V\neq\emptyset$.
Definition 3.3.
Let $X$ and $Y$ be topological spaces and let $\Phi:X\to 2^{Y}$ be a set-valued mapping. A function $f:X\to Y$ is called a selection of $\Phi$, if $f(x)\in\Phi(x)$ for all $x\in X$.
Theorem 3.4 (Modification of Michael’s Selection Theorem (Th. 2 in RS )).
The following properties of a $T_{1}$-space $X$ are equivalent:
(a) $X$ is normal;
(b) The real line $\mathbb{R}$ is an extensor for $X$;
(c) Every separable Banach space is extensor for $X$;
(d) Every lower semicontinuous set-valued mapping $\Phi:X\to\mathcal{C}(\mathbb{R})$ admits a continuous selection;
(e) If $Y$ is a separable Banach space, then every lower semicontinuous set-valued mapping $\Phi:X\to\mathcal{C}(Y)$ admits a continuous selection.
Here $\mathcal{C}(Y)$ denotes the family of all nonempty compact convex subsets of $Y$ or the whole space $Y$.
Now we only have to construct a suitable set-valued mapping.
Definition 3.5.
Suppose we are given a topological space $X$ and a function $f:X\to\mathbb{R}$. For all $x\in X$ we use the following notation:
  • 1.
    $f^{+}:X\to\mathbb{R}$ is the function
    $f^{+}(x)=\inf\limits_{O(x)}\sup\limits_{y\in O(x)}f(x);$
  • 2.
    $f^{-}:X\to\mathbb{R}$ is the function
    $f^{-}(x)=\sup\limits_{O(x)}\inf\limits_{y\in O(x)}f(x);$
  • 3.
    $\Phi_{f}(x)=[f^{+}(x)-\frac{1}{2}\mathrm{dc}(f),f^{-}(x)+\frac{1}{2}\mathrm{dc}(f)]$.
It is easy to see that $f^{+}(x)-f^{-}(x)=\mathrm{dc}(f,x)$ and $\Phi(x)\neq\emptyset$.
Lemma 3.6.
For any topological space $X$ and function $f:X\to\mathbb{R}$ the function $\Phi_{f}:X\to 2^{\mathbb{R}}$ is lower semicontinuous.
Proof.
Let us take any $x_{0}\in X$, any $y_{0}\in\Phi_{f}(x_{0})$ and any $\varepsilon>0$. There is a neighborhood $U$ of the point $x_{0}$ such that for every $x\in U$ we have
$f^{+}(x)<f^{+}(x_{0})+\varepsilon\text{ and }f^{-}(x)>f^{-}(x_{0})-\varepsilon.$
It follows that
$f^{+}(x)-\frac{1}{2}\mathrm{dc}(f)<f^{+}(x_{0})-\frac{1}{2}\mathrm{dc}(f)+\varepsilon\leq y_{0}+\varepsilon$
and
$f^{-}(x)+\frac{1}{2}\mathrm{dc}(f)<f^{-}(x_{0})+\frac{1}{2}\mathrm{dc}(f)-\varepsilon\leq y_{0}-\varepsilon,$
which means that $(y_{0}-\varepsilon,y_{0}+\varepsilon)\cap\Phi_{f}(x)\neq\emptyset$. ∎
Theorem 3.7.
If $X$ is a normal space, then for every function $f:X\to\mathbb{R}$ there is a continuous function $g:X\to\mathbb{R}$ such that
$|f(x)-g(x)|\leq\frac{\mathrm{dc}(f)}{2}$
for all $x\in X$.
Proof.
We apply modification of Michael’s Selection Theorem to $\Phi_{f}$ and obtain some continuous selection $g$. It is easy to see that $|f(x)-g(x)|\leq\frac{1}{2}\mathrm{dc}(f)$ for all $x\in X$. ∎
Remark 3.8.
It is obvious that the upper bound $\frac{1}{2}\mathrm{dc}(f)$ is the best possible.
Now let us denote $\mathbb{H}$ the Hilbert cube $\prod\limits_{n\in\mathbb{N}}[0,\frac{1}{n}]$ with the metric $d_{\mathbb{H}}$ defined as following:
$d_{\mathbb{H}}(x,y)=\max\limits_{n\in\mathbb{N}}|x(n)-y(n)|.$
It is easy to see that the metric $d_{\mathbb{H}}$ is equivalent to the most common euclidean metric of $\mathbb{H}$. The known fact that the space $\mathbb{H}$ is universal for all separable metrizable spaces must explain our attention to the following
Theorem 3.9.
If $X$ is a normal space, then for every function $f:X\to\mathbb{H}$ there is a continuous function $g:X\to\mathbb{H}$ such that
$d_{\mathbb{H}}(f(x),g(x))\leq\frac{\mathrm{dc}(f)}{2}$
for all $x\in X$.
Proof.
Suppose $f(x)=\{f_{n}(x)\}_{n\in\mathbb{N}}$. For every $n\in\mathbb{N}$ choose some continuous function $g_{n}:X\to\mathbb{R}$ such that $|f_{n}(x)-g_{n}(x)|\leq\frac{\mathrm{dc}(f_{n})}{2}\leq\frac{\mathrm{dc}(f)}{2}$ for all $x\in X$. Define $g(x)=\{g_{n}(x)\}_{n\in\mathbb{N}}$ for all $x\in X$. It is easy to see that $g:X\to\mathbb{H}$ is continuous and $d_{\mathbb{H}}(f(x),g(x))\leq\frac{\mathrm{dc}(f)}{2}$. ∎
4 Domain approximation
Definition 4.1.
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. We say that a countable family $\mathcal{E}$ of closed subsets of $X$ is a $\mathrm{DA}$-family (or domain approximation family) for the function $f$ if for every finite $K\subseteq X$ and any $\varepsilon>0$ there is a set $F\in\mathcal{E}$ such that $K\subseteq F$ and $\mathrm{dc}(f|_{F})<\varepsilon$.
Definition 4.2.
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. For all $\varepsilon>0$ we denote $D_{\varepsilon}(f)=\{x\in X:\mathrm{dc}(f,x)\geq\varepsilon\}$. We also denote $D(f)=\bigcup\limits_{\varepsilon>0}D_{\varepsilon}(f)$, which is the set of all discontinueties of $f$.
Lemma 4.3.
Suppose we are given metric spaces $X$ and $Y$ and a function $f:X\to Y$ such that $|D(f)|\leq\omega$. Then there is a $\mathrm{DA}$-family for $f$.
Proof.
Let us denote $I=\mathbb{N}\times\mathbb{N}\times[D(f)]^{<\omega}$. For all $\gamma=(n,k,M)\in I$ we define
$F_{\gamma}=M\cup X\setminus O_{1/k}(D_{1/n}(f)).$
It is obvious that the family $\{F_{\gamma}:\gamma\in I\}$ is a $\mathrm{DA}$-family for $f$. ∎
Corollary 4.4.
If $X$ is a metrizable Luzin space, then for every metric space $Y$ and any $f\in Q(X,Y)$ there is a $\mathrm{DA}$-family for $f$.
5 Flattening lemma
For every set $S$ we denote $(H_{S},\rho_{S})$ the metric $S$-hedgehog space, i.e. $H_{S}=\{0\}\cup((0,1]\times S)$,
$\rho_{S}((p,\alpha),(q,\beta))=\begin{cases}|p-q|,\;\alpha=\beta;\\ p+q,\;\alpha\neq\beta\end{cases}$
and $\rho_{S}(0,(p,\alpha))=x$.
Recall that the $\omega$-th power of the space $H_{S}$ is an universal space for all metrizable spaces of weight $\leq|S|$ (see Eng ). We fix on the space $H_{S}^{\omega}$ the metric
$\rho_{S}^{\omega}(x,y)=\sup\limits_{n\in\omega}\frac{1}{n+1}\rho_{S}(x(n),y(n)),$
which generates the Tychonoff production topology.
Definition 5.1.
Let $S$ be a set and suppose we are given $D\subseteq S$ and $\gamma\in D$. We define the function $h_{S}^{D,\gamma}:H_{S}\to H_{D}$ in the following way: $h_{S}^{D,\gamma}(0)=0$ and
$h_{D}^{\gamma}(x,\alpha)=\begin{cases}(x,\alpha),\;\alpha\in D;\\ (x,\gamma),\;\alpha\notin D.\end{cases}$
Let us notice an obvious property of the function $h_{D}^{\gamma}$.
Proposition 5.2.
Let $S$ be a set and suppose we are given $D\subseteq S$ and $\gamma\in D$. Denote $C=D\setminus\{\gamma\}$. Then for every $x\in H_{C}$ and $y\in H_{S}$ we have $\rho_{S}(h_{D}^{\gamma}(x),h_{D}^{\gamma}(y))=\rho_{S}(x,y)$.
Lemma 5.3.
Suppose we are given a metrizable space $X$ and its separable subset $A$. Then there is a continuous function $\varphi:X\to\mathbb{H}$ such that:
    $\varphi[U\cap A]\subseteq\mathrm{Int}_{\varphi[Y]}f[U]$ (for all open $U\subseteq Y$).
Proof.
We can suppose that $X\subseteq H_{S}^{\omega}$ for some uncountable set $S$ (Eng, , Theorem 4.4.9). Since $A$ is separable, we can choose a countable subset $C\subseteq S$ such that $A\subseteq H_{C}^{\omega}$. Choose any $\gamma\in S\setminus C$ and denote $D=C\cup\{\gamma\}$. Now for all $x\in X$ define $h(x)=\{h_{D}^{\gamma}(x(n))\}_{n\in\omega}$, so $h:X\to H_{D}^{\omega}$. It follows from the previous propostion that $\rho_{S}^{\omega}(h(x),h(y))=\rho_{S}^{\omega}(x,y)$ for all $x\in A$ and $y\in X$. Finally, we take any homeomorphic embedding $g:H_{D}\to\mathbb{H}$ and define $\varphi(x)=g(h(x))$ for all $x\in X$. The function $\varphi$ is as required. ∎
Lemma 5.4.
Let $X$ be a space, $Y$ be a metrizable space and suppose we are given a function $f:X\to Y$ such that its image $A=f[X]$ is separable. Take a function $\varphi:Y\to\mathbb{H}$ with the properties from Lemma 5.3. Then for every family $P$ of functions $X\to Y$ the following conditions are equivalent:
  1. (1)
    $f\in\overline{P}$;
  2. (2)
    $f\circ\varphi\in\overline{\{h\circ\varphi:h\in P\}}$.
Proof.
$(1)\to(2)$ follows from continuity of $\varphi$.
$(2)\to(1)$ follows from the property (*). ∎
6 Application to types of tightness of spaces of quasicontinuous functions
Definition 6.1.
Suppose we are given a topological space $X$, a metric space $(Y,d)$, a function $f:X\to Y$ and a familly $P\subseteq Y^{X}$.
  • 1.
    For every $r\geq 0$ we write $d(f,P)\leq r$ if for every finite $K\subseteq X$ there is a function $h\in P$ such that $d(f(x),h(x))\leq r$ for all $x\in K$.
  • 2.
    We denote $\textbf{d}(f,P)=\inf\{r\geq 0:d(f,P)\leq r\}$.
It is obvious that $f\in\overline{P}$ if and only if $\textbf{d}(f,P)=0$.
6.1 Tightness
Lemma 6.2.
Let $X$ be an open Whyburn topological space and $(Y,d)$ be metric space, $f\in C(X,Y)$ and suppose that $X$ is $\mathcal{K}_{\Omega}$-Lindelöf. Then for every $P\subseteq Q(X,Y)$ such that $\textbf{d}(f,P)\leq\varepsilon$ there is a subset $S\in[P]^{\leq\omega}$ such that $\textbf{d}(f,S)\leq\varepsilon$.
Proof.
For every $h\in P$ and $n\in\mathbb{N}$ denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in Lev ).
Since $\textbf{d}(f,P)\leq\varepsilon$, $\mathcal{V}_{n}=\{V_{h,n}:h\in P\}$ is a semi-open $\omega$-cover of $X$ for every $n\in\mathbb{N}$.
Since $X$ is an open Whyburn topological space, for every $V_{h,n}$ and a finite subset $K$ of $V_{h,n}$ we can consider an open set $W_{K,h,n}$ such that $K\subset\overline{W_{K,h,n}}\subset V_{h,n}$. Then the family $\{W_{K,h,n}:K\in[V_{h,n}]^{<\omega}$ and $h\in P\}\in\mathcal{K}_{\Omega}$ for every $n\in\mathbb{N}$. For each $n\in\mathbb{N}$ we apply that $X$ is $\mathcal{K}_{\Omega}$-Lindelöf and we get the countable family $\{W_{K_{i}(n),h_{i}(n),n}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$. It remains to note that $S=\{h_{i}(n):i,n\in\mathbb{N}\}$ such that $S\in[P]^{\leq\omega}$ and $\textbf{d}(f,S)\leq\varepsilon$.
Lemma 6.3.
Let $X$ be a metrizable $\mathcal{K}_{\Omega}$-Lindelöf space and $F$ be a closed subset of $X$. Then $F$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Proof.
(1) Assume that $IntF=\emptyset$. Since $X$ is Lusin, $F$ is countable and, hence, $F$ is $\mathcal{K}_{\Omega}$-Lindelöf.
(2) Let $IntF\neq\emptyset$. By Theorem 3.6 ($(2)\Rightarrow(5)(a)$) in Osipov (Any open subset of $\mathcal{K}_{\Omega}$-Lindelöf space is $\mathcal{K}_{\Omega}$-Lindelöf), $IntF$ is $\mathcal{K}_{\Omega}$-Lindelöf and, by Theorem 3.6 ($(2)\Rightarrow(5)(b)$) (The union of $\mathcal{K}_{\Omega}$-Lindelöf space with countable space is $\mathcal{K}_{\Omega}$-Lindelöf), $F=IntF\cup(F\setminus IntF)$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Theorem 6.4.
For every metrizable space $X$ the following conditions are equivalent:
  1. 1.
    $t(Q_{p}(X,Y))=\omega$ for all metrizable spaces $Y$;
  2. 2.
    There is a metrizable space $Y$ and a function $f\in Q_{p}(X,Y)$ such that $t(f,Q_{p}(X,Y))=\omega$;
  3. 3.
    There is a function $f\in Q_{p}(X,\mathbb{R})$ such that $t(f,Q_{p}(X,\mathbb{R}))=\omega$;
  4. 4.
    $X$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Proof.
$(1)\Rightarrow(2)$ and $(1)\Rightarrow(3)$ are trivial.
$(2)\Rightarrow(4)$. (i) Suppose that $t(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)\neq Y$ where $Y$ is metrizable and $|Y|>1$. Let a family $\gamma=\{V_{\alpha}\}$ of open subsets of $X$ such that $\gamma\in\mathcal{K}_{\Omega}$. Let $y\in Y\setminus f(X)$.
For every $K\in[X]^{<\omega}$ there is $V_{\alpha}\in\gamma$ such that $K=\{x_{1},...,x_{k}\}\subset\overline{V_{\alpha}}$. Since $X$ is metrizable (hence, it is open Whyburn), there are open subsets $W_{1}$,…,$W_{k}$ of $X$ such that $x_{i}\in\overline{W_{i}}\subset V_{\alpha}\cup\{x_{i}\}$ for $i=1,...,k$, $\overline{W_{i}}\cap\overline{W_{j}}=\emptyset$ for $i\neq j$. Consider the function
$f_{K,\alpha}:=\left\{\begin{array}[]{lcl}f(x_{i})\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\overline{W_{i}};\\ y\,\,\,\,\,\,\,\,on\,\,\,X\setminus(\bigcup\{\overline{W_{i}}:i=1,...,k\}).\\ \end{array}\right.$
Let $F=\{f_{K,\alpha}:K\in[X]^{<\omega}\}$. It is clear that $f\in\overline{F}$. Then there is a countably subset $S=\{f_{K_{i},\alpha_{i}}:i\in\mathbb{N}\}$ of $F$ such that $f\in\overline{S}$.
Claim that $\{V_{\alpha_{i}}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$. Let $K\in[X]^{<\omega}$. Consider the open neighborhood $[K,Y\setminus\{y\}]:=\{h\in Q_{p}(X,Y):h(x)\in Y\setminus\{y\}$ for every $x\in K\}$ of the point $f$ in the space $Q_{p}(X,Y)$. Then, there is $f_{K_{i},\alpha_{i}}\in S\cap[K,Y\setminus\{y\}]$. It follows that $K\subset\overline{V_{\alpha_{i}}}$ and, hence, $\{V_{\alpha_{i}}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
(ii) Suppose that $t(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)=Y$ where $Y$ is a metrizable space with a metric $d$.
(1) Claim that there are non-empty open disjoint subsets $U$ and $V$ of $Y$ such that $\overline{Intf^{-1}(U)\cup Intf^{-1}(V)}=X$.
Choose $p,q\in Y$ such that $p\neq q$ and for each $\varepsilon\in(0,d(p,q))$ we consider $U_{\varepsilon}=O_{\varepsilon}(p)$ and $V_{\varepsilon}=X\setminus\overline{O_{\varepsilon}(p)}$. It is clear that $U_{\varepsilon}\cap V_{\varepsilon}=\emptyset$. Note that if $\overline{Intf^{-1}(U_{\varepsilon})\cup Intf^{-1}(V_{\varepsilon})}\neq X$ then $Int(f^{-1}[FrO_{\varepsilon}(p)])\neq\emptyset$. Let us show that for all $\varepsilon$ such a situation cannot exist.
Let’s assume the opposite. For each $\varepsilon\in(0,d(p,q))$ we choose a point $x_{\varepsilon}\in Int(f^{-1}[FrO_{\varepsilon}(p)])$ and a neighborhood $W_{\varepsilon}$ of $x_{\varepsilon}$ such that $\overline{W_{\varepsilon}}\subset Int(f^{-1}[FrO_{\varepsilon}(p)])$. For each $E\subset(0,d(p,q))$ we define the function $h_{E}$:
$h_{E}:=\left\{\begin{array}[]{lcl}p\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\bigcup\{\overline{W_{\varepsilon}}:\varepsilon\in E\};\\ f\,\,\,\,\,\,\,\,on\,\,\,X\setminus(\bigcup\{\overline{W_{\varepsilon}}:\varepsilon\in E\}).\\ \end{array}\right.$
Let $Z=\{h_{E}:E\subset(0,d(p,q))\}$. Note that $Z\subset Q(X,Y)$ and $f=h_{\emptyset}\in Z$. It is clear that $Z$ is homeomorphic to $2^{\mathfrak{c}}$ and, hence, $t(f,Q_{p}(X,Y))>\omega$.
(2) Let $R=Y\setminus(U\cup V)$ where $U$ and $V$ from (1). For each $x\in f^{-1}(R)$ we choose $W_{x}\in\{U,V\}$ such that $f\upharpoonright(f^{-1}[W_{x}]\cup\{x\})$ is quasicontinuous. Let $A=f^{-1}(U)\cup\{x\in f^{-1}(R):W_{x}=U\}$ and $B=f^{-1}(V)\cup\{x\in f^{-1}(R):W_{x}=V\}$. Thus, $X=A\sqcup B$ and $f\upharpoonright A$ and $f\upharpoonright B$ are quasicontinuous.
Let $p\in U$ and $q\in V$.
(3) Let a family $\gamma=\{W_{\alpha}\}$ of open subsets of $X$ such that $\gamma\in\mathcal{K}_{\Omega}$.
We can represent each set $W_{\alpha}$ as a finite disjoint union $W^{1}_{\alpha}\sqcup...\sqcup W^{n}_{\alpha}$ of open sets $W^{j}_{\alpha}$ where $W^{j}_{\alpha}\subset Intf^{-1}(U)$ or $W^{j}_{\alpha}\subset Intf^{-1}(V)$ for each $j=1,...,n$ and $|\overline{W^{j}_{\alpha}}\cap FrA|\leq 1$. Note that since $X$ is an open Whyburn space, we can assume this property.
For each $W_{\alpha}=W^{1}_{\alpha}\sqcup...\sqcup W^{n}_{\alpha}\in\gamma$ and a finite subset $K=\{x_{1},...,x_{n}\}$ of $X$ where $x_{j}\in\overline{W^{j}_{\alpha}}$, $j=1,...,n$, we define the function $h_{W_{\alpha},K}$:
$h_{W_{\alpha},K}:=\left\{\begin{array}[]{lcl}f(x_{j})\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\overline{W^{j}_{\alpha}};\\ q\,\,\,\,\,\,\,\,on\,\,\,A\setminus W_{\alpha}\\ p\,\,\,\,\,\,\,\,on\,\,\,B\setminus W_{\alpha}.\\ \end{array}\right.$
Note that $h_{W_{\alpha},K}\in Q(X,Y)$ and $f\in\overline{\{h_{W_{\alpha},K}:W_{\alpha}\in\gamma\}}$.
Since $t(f,Q_{p}(X,Y))\leq\omega$, there is a countable family $\{h_{W_{\alpha_{i}},K_{i}}:i\in\mathbb{N}\}$ such that $f\in\overline{\{h_{W_{\alpha_{i}},K_{i}}:i\in\mathbb{N}\}}$. It is remain that $\{W_{\alpha_{i}}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
$(4)\Rightarrow(1)$. By Lemma 3.1 in Osipov , $X$ is a Lusin space.
Let $f\in Q_{p}(X,Y)$ and $f\in\overline{P}$ where $P\subset Q_{p}(X,Y)$.
Let us first reduce the general case to the case $Y=\mathbb{H}$. Since $f$ is a quasicontinuous function on a Lusin space, the set $A=f[X]$ is separable. Thus, by Lemma 5.4 there is a function $\varphi:Y\to\mathbb{H}$ such that for every family $T$ of functions $X\to Y$ we have $f\in\overline{T}$ if and only if $f\circ\varphi\in\overline{\{h\circ\varphi:h\in T\}}$, so now we only have to consider the case $Y=\mathbb{H}$.
By Corollary 4.4, there is a $\mathrm{DA}$-family $\mathcal{E}$ for $f$. For every $F\in\mathcal{E}$ denote by $\varepsilon_{F}:=\mathrm{dc}(f,F)$. Thus, for every $F\in\mathcal{E}$ there is $g_{F}\in C(F,Y)$ such that $d(f(x),g_{F}(x))\leq\varepsilon_{F}$ for every $x\in F$.
By Lemma 6.3, $F$ is $\mathcal{K}_{\Omega}$-Lindelöf. Note that $d(g_{F},P\upharpoonright F)\leq\varepsilon_{F}$. By Lemma 6.5, there is $T_{F}\in[P]^{\leq\omega}$ such that $d(g_{F},T\upharpoonright F)\leq\varepsilon_{F}$.
Consider $T=\bigcup\{T_{F}:F\in\mathcal{E}\}$. It is remain note that $T\in[P]^{\leq\omega}$ and $f\in\overline{T}$.
$(3)\Rightarrow(4)$. It is the implication $(2)\Rightarrow(4)$ for $Y=\mathbb{R}$.
6.2 Fan-tightness
Lemma 6.5.
Let $X$ be an open Whyburn topological space and $(Y,d)$ be metric space, $f\in C(X,Y)$ and suppose that $X$ satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$. Then for every $P_{n}\subseteq Q(X,Y)$ such that $\textbf{d}(f,P_{n})\leq\varepsilon$ there is a finite subset $S_{n}$ of $P_{n}$ such that $\textbf{d}(f,\bigcup S_{n})\leq\varepsilon$.
Proof.
For every $n\in\mathbb{N}$ and $h\in P_{n}$ and denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in Lev ).
Since $\textbf{d}(f,P)\leq\varepsilon$, $\mathcal{V}_{n}=\{V_{h,n}:h\in P\}$ is a semi-open $\omega$-cover of $X$ for every $n\in\mathbb{N}$.
Since $X$ is an open Whyburn topological space, for every $V_{h,n}$ and a finite subset $K$ of $V_{h,n}$ we can consider an open set $W_{K,h,n}$ such that $K\subset\overline{W_{K,h,n}}\subset V_{h,n}$. Then the family $\{W_{K,h,n}:K\in[V_{h,n}]^{<\omega}$ and $h\in P_{n}\}\in\mathcal{K}_{\Omega}$ for every $n\in\mathbb{N}$.
Since $X$ satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$, for each $n\in\mathbb{N}$ there is a finite family $T_{n}=\{W_{K_{1},h_{1},n},...,W_{K_{s(n)},h_{s(n)}}\}$ such that $\bigcup\{T_{n}:n\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
It remains to note that the family $\{S_{n}:S_{n}=\{h_{1},...,h_{s(n)}\}\subset P_{n},n\in\mathbb{N}\}$ such that $\textbf{d}(f,\bigcup S_{n})\leq\varepsilon$.
Theorem 6.6.
For every metrizable space $X$ the following conditions are equivalent:
  1. 1.
    $vet(Q_{p}(X,Y))=\omega$ for all metric spaces $Y$;
  2. 2.
    There is a metric space $Y$ and a function $f\in Q_{p}(X,Y)$ such that $vet(f,Q_{p}(X,Y))=\omega$;
  3. 3.
    There is a function $f\in Q_{p}(X,\mathbb{R})$ such that $vet(f,Q_{p}(X,\mathbb{R}))=\omega$;
  4. 4.
    $X$ satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$.
Proof.
$(1)\Rightarrow(2)$ and $(1)\Rightarrow(3)$ are trivial.
$(2)\Rightarrow(4)$. (i) Suppose that $vet(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)\neq Y$ where $Y$ is metrizable and $|Y|>1$.
Let $\gamma_{i}=\{V^{i}_{\alpha}\}$ be a family of open subsets of $X$ such that $\gamma_{i}\in\mathcal{K}_{\Omega}$ for each $i\in\mathbb{N}$ and $y\in Y\setminus f(X)$.
For every $i\in\mathbb{N}$ and $K\in[X]^{<\omega}$ there is $V^{i}_{\alpha}\in\gamma_{i}$ such that $K=\{x_{1},...,x_{k}\}\subset\overline{V^{i}_{\alpha}}$. Since $X$ is metrizable (hence, it is open Whyburn), there are open subsets $W_{1}$,…,$W_{k}$ of $X$ such that $x_{s}\in\overline{W_{s}}\subset\overline{V^{i}_{\alpha}}$ for $s=1,...,k$, $\overline{W_{m}}\cap\overline{W_{j}}=\emptyset$ for $m\neq j$. Consider the function
$f^{i}_{K,\alpha}:=\left\{\begin{array}[]{lcl}f(x_{s})\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\overline{W_{s}};\\ y\,\,\,\,\,\,\,\,on\,\,\,X\setminus(\bigcup\{\overline{W_{s}}:s=1,...,k\}).\\ \end{array}\right.$
Let $F_{i}=\{f^{i}_{K,\alpha}:K\in[X]^{<\omega}\}$. It is clear that $f\in\overline{F_{i}}$. Then for every $i\in\mathbb{N}$ there is a finite subset $S_{i}=\{f^{i}_{K_{1},\alpha_{1}},...,f^{i}_{K_{l(i)},\alpha_{l(i)}}\}$ of $F_{i}$ such that $f\in\overline{\bigcup\{S_{i}:i\in\mathbb{N}\}}$.
Claim that $\{V^{i}_{\alpha_{p}}:p\in\{1,...,l(i)\},i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
Let $K\in[X]^{<\omega}$. Consider the open neighborhood $[K,Y\setminus\{y\}]:=\{h\in Q_{p}(X,Y):h(x)\in Y\setminus\{y\}$ for every $x\in K\}$ of the point $f$ in the space $Q_{p}(X,Y)$. Then, there is $i\in\mathbb{N}$ and $p\in\{1,...,l(i)\}$ $f^{i}_{K_{p},\alpha_{p}}\in S\cap[K,Y\setminus\{y\}]$. It follows that $K\subset\overline{V^{i}_{\alpha_{p}}}$ and, hence, $\{V^{i}_{\alpha_{p}}:p\in\{1,...,l(i)\},i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
(ii) If $vet(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)=Y$ where $Y$ is metrizable and $|Y|>1$. Let $\gamma_{i}=\{W^{i}_{\alpha}\}$ be a family of open subsets of $X$ such that $\gamma_{i}\in\mathcal{K}_{\Omega}$ for each $i\in\mathbb{N}$. Then, similar to implication $(2)\Rightarrow(4)$ in Theorem 6.4, we construct the function $h_{W_{\alpha,i},K}$ for each $W_{\alpha,i}\in\gamma_{i}$.
Let $F_{i}=\{h_{W_{\alpha,i},K}:K\in[X]^{<\omega}\}$. It is clear that $f\in\overline{F_{i}}$. Then for every $i\in\mathbb{N}$ there is a finite subset $S_{i}=\{h^{1}_{W_{\alpha_{1},i},K_{1}},...,h^{n}_{W_{\alpha_{n(i)},i},K_{n(i)}}\}$ of $F_{i}$ such that $f\in\overline{\bigcup\{S_{i}:i\in\mathbb{N}\}}$. It is remain note that $\{W_{\alpha_{p},i}:p\in\{1,...,n(i)\},i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
$(3)\Rightarrow(1)$. It is clear that any space satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$ is $\mathcal{K}_{\Omega}$-Lindelöf. Then, by Lemma 3.1 in Osipov , $X$ is a Lusin space.
Let $f\in Q_{p}(X,Y)$ and $f\in\overline{P_{n}}$ where $P_{n}\subset Q_{p}(X,Y)$ for each $n\in\mathbb{N}$.
Similar to the previous theorem, now we only have to consider the case $Y=\mathbb{H}$.
By Corollary 4.4, there is a $\mathrm{DA}$-family $\mathcal{E}$ for $f$. For every $F\in\mathcal{E}$ denote by $\varepsilon_{F}:=\mathrm{dc}(f,F)$. Thus, for every $F\in\mathcal{E}$ there is $g_{F}\in C(F,Y)$ such that $d(f(x),g_{F}(x))\leq\varepsilon_{F}$ for every $x\in F$.
By Lemma 6.3, $F$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Note that $d(g_{F},P_{n}\upharpoonright F)\leq\varepsilon_{F}$. By Lemma 6.5, there is $T_{F}\in[P_{n}]^{<\omega}$ such that $d(g_{F},T_{F}\upharpoonright F)\leq\varepsilon_{F}$.
Consider $T=\bigcup\{T_{F}:F\in\mathcal{E}\}$. It is remain note that $f\in\overline{T}$.
$(3)\Rightarrow(4)$. It is the implication $(2)\Rightarrow(4)$ for $Y=\mathbb{R}$.
6.3 Strong fan-tightness
Similar to the previous theorem, the following result is easy to prove.
Theorem 6.7.
For every metrizable space $X$ the following conditions are equivalent:
  1. 1.
    $vet_{1}(Q_{p}(X,Y))=\omega$ for all metric spaces $Y$;
  2. 2.
    There is a metric space $Y$ and a function $f\in Q_{p}(X,Y)$ such that $vet_{1}(f,Q_{p}(X,Y))=\omega$;
  3. 3.
    There is a function $f\in Q_{p}(X,\mathbb{R})$ such that $vet_{1}(f,Q_{p}(X,\mathbb{R}))=\omega$;
  4. 4.
    $X$ satisfies $S_{1}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$.
Combining the results in Os1 ; Osipov and this paper, we get the following diagram in the class of metrizable spaces $X$ and $Y$.
$X$ $Q_{p}(X,\,Y)$
countable Fréchet - Urysohn
$S_{1}(K_{\Omega},\,K_{\Omega})$ $\mbox{vet}_{1}(Q_{p}(X,\,Y))=\omega$
$S_{fin}(K_{\Omega},\,K_{\Omega})$ $\mbox{vet}(Q_{p}(X,\,Y))=\omega$
$K_{\Omega}$-Lindelöf $t(Q_{p}(X,\,Y))=\omega$
Diagram 1.
Remark 6.8.
In (Osipov , Example 6.1) it was shown that, under Jensen’s axiom, there exists a metrizable Lusin space $X$ such that $vet_{1}(Q_{p}(X,\mathbb{R}))=\omega$. An interesting open question (Questions 2 and 3 in Osipov ) remains: Is there a $T_{2}$-space $X$ such that $Q_{p}(X,\mathbb{R})$ has countable tightness but none $Q_{p}(X,\mathbb{R})$ has countable (strong) fan-tightness?