### v2

# Tightness type properties of spaces of quasicontinuous functions

Abstract
Using approximation by continuous functions we prove the following statements to types of tightness in a space $Q_{p}(X,\mathbb{R})$ of all quasicontinuous real-valued functions with the topology $\tau_{p}$ of pointwise convergence: the countability of tightness (fan-tightness, strong fan-tightness) at a point $f$ of space $Q_{p}(X,\mathbb{R})$ implies the countability of tightness (fan-tightness, strong fan-tightness) of space $Q_{p}(X,Y)$ of all quasicontinuous functions from $X$ into any non-one-point metrizable space $Y$. This result is the answer to the open question in the class of metrizable spaces.
keywords:
quasicontinuous function , Lusin space , open Whyburn space , tightness , fan-tightness , strong fan-tightness , selection principle , continuous selection
MSC:
[2020] 54C35 , 54A25 , 54D20 , 54C10
journal:
1 Introduction
A function $f:X\rightarrow Y$ is quasicontinuous at $x$ if for any open set $V$ containing $f(x)$ and any $U$ open containing $x$, there exists a nonempty open set $W\subseteq U$ such that $f(W)\subseteq V$. It is quasicontinuous if it is quasicontinuous at every point. Call a set semi-open (or quasi-open) if it is contained in the closure of its interior. Then $f:X\rightarrow Y$ is quasicontinuous if and only if the inverse of every open set is quasi-open.
Let $X$ and $Y$ be Hausdorff topological spaces, $Q_{p}(X,Y)=(Q(X,Y),\tau_{p})$ be the space of all quasicontinuous functions on $X$ with values in $Y$ and $\tau_{p}$ be the pointwise convergence topology.
We were motivated for this paper by the problem of tightness in a space of quasi-continuous functions.
It is an open problem whether the space $Q_{p}(X,\mathbb{R})$ is topologically homogeneous (note that this space is not a linear space, since the sum of quasi-continuous functions is not necessarily quasi-continuous). In particular, it is unknown whether the points of $Q_{p}(X,\mathbb{R})$ can be distinguished using some topological property or cardinal invariant. In this regard, in ([5], Question 1) the second author raised a question on tightness in the space $Q_{p}(X,\mathbb{R})$. It was proved in [5] that for every continuous $f:X\to\mathbb{R}$ we have $t(f,Q_{p}(X,\mathbb{R}))=\omega$ if and only if the space $X$ is $\mathcal{K}_{\Omega}$-Lindelof, and it was unknown whether the same is true for non-continuous functions.
In this paper we prove that the property of countable tightness (as well as countable fan-tightness and countable strong fan-tightness) can not distinguish points of the space $Q_{p}(X,Y)$ whenever $Y$ is metrizable.
One of the main tools of our proof is Theorem 3.1 below on approximation of arbitrary functions by continuous functions (in the metric of uniform convergence).
2 Preliminaries
Let us recall some properties of a topological space $X$.
(1) A space $X$ has countable tightness at a point $x$ (denoted $t(x,X)=\omega$) if $x\in\overline{A}$, then $x\in\overline{B}$ for some countable $B\subseteq A$. A space $X$ has countable tightness (denoted $t(X)=\omega$) if $t(x,X)=\omega$ for every $x\in X$.
(2) A space $X$ has countable fan-tightness at a point $x$ (denoted $vet(x,X)=\omega$) if for any countable family $\{A_{n}:n\in\omega\}$ of subsets of $X$ satisfying $x\in\bigcap_{n\in\omega}\overline{A_{n}}$ it is possible to select finite sets $K_{n}\subset A_{n}$ in such a way that $x\in\overline{\bigcup_{n\in\omega}K_{n}}$. A space $X$ has countable fan-tightness (denoted $vet(X)=\omega$) if $vet(x,X)=\omega$ for every $x\in X$.
(3) A space $X$ is said to have countable strong fan-tightness at a point $x$ (denoted $vet_{1}(x,X)=\omega$) if for each countable family $\{A_{n}:n\in\omega\}$ of subsets of $X$ such that $x\in\bigcap_{n\in\omega}\overline{A_{n}}$, there exist $a_{i}\in A_{i}$ such that $x\in\overline{\{a_{i}:i\in\omega\}}$. A space $X$ has countable strong fan-tightness (denoted $vet_{1}(X)=\omega$) if $vet_{1}(x,X)=\omega$ for every $x\in X$.
(4) A space $X$ is said to be open Whyburn if for every open set $A\subset X$ and every $x\in\overline{A}\setminus A$ there is an open set $B\subseteq A$ such that $\overline{B}\setminus A=\{x\}$ [4].
Note that the class of open Whyburn spaces is quite wide; for example, it includes all first countable regular spaces [4] and, therefore, all metrizable spaces.
Let $X$ be a Tychonoff topological space, $C(X,\mathbb{R})$ be the space of all continuous functions on $X$ with values in $\mathbb{R}$ and $\tau_{p}$ be the pointwise convergence topology. Denote by $C_{p}(X,\mathbb{R})$ the topological space $(C(X,\mathbb{R}),\tau_{p})$.
The symbol $\bf{0}$ stands for the constant function to 00. A basic open neighborhood of $\bf{0}$ in $\mathbb{R}^{X}$ is of the form $[F,(-\varepsilon,\varepsilon)]=\{f\in\mathbb{R}^{X}:f(F)\subset(-\varepsilon,\varepsilon)\}$, where $F\in[X]^{<\omega}$ and $\varepsilon>0$.
Let us recall that a cover $\mathcal{U}$ of a set $X$ is called
$\bullet$ an $\omega$-cover if each finite set $F\subseteq X$ is contained in some $U\in\mathcal{U}$;
$\bullet$ a $\gamma$-cover if for any $x\in X$ the set $\{U\in\mathcal{U}:x\not\in U\}$ is finite.
In this paper $\mathcal{A}$ and $\mathcal{B}$ will be collections of the following covers of a space $X$:
$\mathcal{O}^{s}$ : the collection of all semi-open covers of $X$.
$\Omega$ : the collection of open $\omega$-covers of $X$.
$\mathcal{K}$: the collection $\mathcal{U}$ of open subsets of $X$ such that $X=\bigcup\{\overline{U}:U\in\mathcal{U}\}$.
$\mathcal{K}_{\Omega}$ is the set of $\mathcal{U}$ in $\mathcal{K}$ such that no element of $\mathcal{U}$ is dense in $X$, and for each finite set $F\subseteq X$, there is a $U\in\mathcal{U}$ such that $F\subseteq\overline{U}$.
A space is $\mathcal{K}_{\Omega}$-Lindelöf if each element of $\mathcal{K}_{\Omega}$ has a countable subset in $\mathcal{K}_{\Omega}$ [9].
Definition 2.1.
([3]) A Hausdorff space $X$ is called a Lusin space if
(a) Every nowhere dense set in $X$ is countable;
(b) $X$ has at most countably many isolated points;
(c) $X$ is uncountable.
If $X$ is an uncountable Hausdorff space then $X$ is $\mathcal{O}^{s}$-Lindelöf (semi-Lindelöf) if and only if $X$ is a Lusin space (Corollary 2.5 in [10]).
If $X$ is a Lusin space, $X$ is hereditarily Lindelöf (Lemma 1.2 in [3]). Hence, if $X$ is a regular Lusin space then $X$ is perfect normal (3.8.A.(b) in [2]).
If $X$ is a Lusin space, so is every uncountable subspace (Lemma 1.1 in [3]).
Many topological properties are defined or characterized in terms of the following classical selection principles (see [8]). Let $\mathcal{A}$ and $\mathcal{B}$ be sets consisting of families of subsets of an infinite set $X$. Then:
$S_{1}(\mathcal{A},\mathcal{B})$ is the selection hypothesis: for each sequence $\{A_{n}:n\in\mathbb{N}\}$ of elements of $\mathcal{A}$ there is a sequence $\{b_{n}\}_{n\in\mathbb{N}}$ such that for each $n$, $b_{n}\in A_{n}$, and $\{b_{n}:n\in\mathbb{N}\}\in\mathcal{B}$.
$S_{fin}(\mathcal{A},\mathcal{B})$ is the selection hypothesis: for each sequence $\{A_{n}:n\in\mathbb{N}\}$ of elements of $\mathcal{A}$ there is a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of finite sets such that for each $n$, $B_{n}\subseteq A_{n}$, and $\bigcup_{n\in\mathbb{N}}B_{n}\in\mathcal{B}$.
In this paper, by metric space we mean a non-trivial metric space, that is, of cardinality greater than 1.
For other notation and terminology almost without exceptions we follow the Engelking’s book [2] and the papers [4, 5].
3 Approximation by continuous functions
Theorem 3.1.
Let $X$ be a normal space. For any $f\in\mathbb{R}^{X}$ there exists a continuous function $g:X\rightarrow\mathbb{R}$ such that $|f(x)-g(x)|\leq\frac{osc(f)}{2}$, where $osc(f)$ is the oscillation of $f$.
This result can be found, for instance, in [1] (Proposition 1.18) (see also [2], Problem 1.7.15 (b)).
Now let us denote $\mathbb{H}$ the Hilbert cube $\prod\limits_{n\in\mathbb{N}}[0,\frac{1}{n}]$ with the metric $d_{\mathbb{H}}$ defined as following:
$d_{\mathbb{H}}(x,y)=\max\limits_{n\in\mathbb{N}}|x(n)-y(n)|.$
It is easy to see that the metric $d_{\mathbb{H}}$ is equivalent to the most common euclidean metric of $\mathbb{H}$. The known fact that the space $\mathbb{H}$ is universal for all separable metrizable spaces must explain our attention to the following
Proposition 3.2.
If $X$ is a normal space, then for every function $f:X\to\mathbb{H}$ there is a continuous function $g:X\to\mathbb{H}$ such that
$d_{\mathbb{H}}(f(x),g(x))\leq\frac{osc(f)}{2}$
for all $x\in X$.
Proof.
Suppose $f(x)=\{f_{n}(x)\}_{n\in\mathbb{N}}$. For every $n\in\mathbb{N}$ choose some continuous function $g_{n}:X\to\mathbb{R}$ such that $|f_{n}(x)-g_{n}(x)|\leq\frac{osc(f_{n})}{2}\leq\frac{osc(f)}{2}$ for all $x\in X$. Define $g(x)=\{g_{n}(x)\}_{n\in\mathbb{N}}$ for all $x\in X$. It is easy to see that $g:X\to\mathbb{H}$ is continuous and $d_{\mathbb{H}}(f(x),g(x))\leq\frac{osc(f)}{2}$. ∎
4 Domain approximation
Definition 4.1.
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. We say that a countable family $\mathcal{E}$ of closed subsets of $X$ is a $\mathrm{DA}$-family (or domain approximation family) for the function $f$ if for every finite $K\subseteq X$ and any $\varepsilon>0$ there is a set $F\in\mathcal{E}$ such that $K\subseteq F$ and $osc(f|_{F})<\varepsilon$.
Definition 4.2.
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. For all $\varepsilon>0$ we denote $D_{\varepsilon}(f)=\{x\in X:osc(f,x)\geq\varepsilon\}$. We also denote $D(f)=\bigcup\limits_{\varepsilon>0}D_{\varepsilon}(f)$, which is the set of all discontinueties of $f$.
Lemma 4.3.
Suppose we are given metric spaces $X$ and $Y$ and a function $f:X\to Y$ such that $|D(f)|\leq\omega$. Then there is a $\mathrm{DA}$-family for $f$.
Proof.
Let us denote $I=\mathbb{N}\times\mathbb{N}\times[D(f)]^{<\omega}$. For all $\gamma=(n,k,M)\in I$ we define
$F_{\gamma}=M\cup X\setminus O_{1/k}(D_{1/n}(f)).$
It is obvious that the family $\{F_{\gamma}:\gamma\in I\}$ is a $\mathrm{DA}$-family for $f$. ∎
Corollary 4.4.
If $X$ is a metrizable Lusin space, then for every metric space $Y$ and any $f\in Q(X,Y)$ there is a $\mathrm{DA}$-family for $f$.
5 Flattening lemma
For every set $S$ we denote $(H_{S},\rho_{S})$ the metric $S$-hedgehog space, i.e. $H_{S}=\{0\}\cup((0,1]\times S)$,
$\rho_{S}((p,\alpha),(q,\beta))=\begin{cases}|p-q|,\;\alpha=\beta;\\ p+q,\;\alpha\neq\beta\end{cases}$
and $\rho_{S}(0,(p,\alpha))=x$.
Recall that the $\omega$-th power of the space $H_{S}$ is an universal space for all metrizable spaces of weight $\leq|S|$ (see [2]). We fix on the space $H_{S}^{\omega}$ the metric
$\rho_{S}^{\omega}(x,y)=\sup\limits_{n\in\omega}\frac{1}{n+1}\rho_{S}(x(n),y(n)),$
which generates the Tychonoff production topology.
Definition 5.1.
Let $S$ be a set and suppose we are given $D\subseteq S$ and $\gamma\in D$. We define the function $h_{S}^{D,\gamma}:H_{S}\to H_{D}$ in the following way: $h_{S}^{D,\gamma}(0)=0$ and
$h_{D}^{\gamma}(x,\alpha)=\begin{cases}(x,\alpha),\;\alpha\in D;\\ (x,\gamma),\;\alpha\notin D.\end{cases}$
Let us notice an obvious property of the function $h_{D}^{\gamma}$.
Proposition 5.2.
Let $S$ be a set and suppose we are given $D\subseteq S$ and $\gamma\in D$. Denote $C=D\setminus\{\gamma\}$. Then for every $x\in H_{C}$ and $y\in H_{S}$ we have $\rho_{S}(h_{D}^{\gamma}(x),h_{D}^{\gamma}(y))=\rho_{S}(x,y)$.
Lemma 5.3.
Suppose we are given a metrizable space $X$ and its separable subset $A$. Then there is a continuous function $\varphi:X\to\mathbb{H}$ such that:
    $\varphi[U\cap A]\subseteq\mathrm{Int}_{\varphi[Y]}f[U]$ (for all open $U\subseteq Y$).
Proof.
We can suppose that $X\subseteq H_{S}^{\omega}$ for some uncountable set $S$ [2, Theorem 4.4.9]. Since $A$ is separable, we can choose a countable subset $C\subseteq S$ such that $A\subseteq H_{C}^{\omega}$. Choose any $\gamma\in S\setminus C$ and denote $D=C\cup\{\gamma\}$. Now for all $x\in X$ define $h(x)=\{h_{D}^{\gamma}(x(n))\}_{n\in\omega}$, so $h:X\to H_{D}^{\omega}$. It follows from the previous propostion that $\rho_{S}^{\omega}(h(x),h(y))=\rho_{S}^{\omega}(x,y)$ for all $x\in A$ and $y\in X$. Finally, we take any homeomorphic embedding $g:H_{D}\to\mathbb{H}$ and define $\varphi(x)=g(h(x))$ for all $x\in X$. The function $\varphi$ is as required. ∎
Lemma 5.4.
Let $X$ be a space, $Y$ be a metrizable space and suppose we are given a function $f:X\to Y$ such that its image $A=f[X]$ is separable. Take a function $\varphi:Y\to\mathbb{H}$ with the properties from Lemma 5.3. Then for every family $P$ of functions $X\to Y$ the following conditions are equivalent:
  1. (1)
    $f\in\overline{P}$;
  2. (2)
    $f\circ\varphi\in\overline{\{h\circ\varphi:h\in P\}}$.
Proof.
$(1)\to(2)$ follows from continuity of $\varphi$.
$(2)\to(1)$ follows from the property (*). ∎
6 Application to types of tightness of spaces of quasicontinuous functions
Definition 6.1.
Suppose we are given a topological space $X$, a metric space $(Y,d)$, a function $f:X\to Y$ and a familly $P\subseteq Y^{X}$.
  • 1.
    For every $r\geq 0$ we write $d(f,P)\leq r$ if for every finite $K\subseteq X$ there is a function $h\in P$ such that $d(f(x),h(x))\leq r$ for all $x\in K$.
  • 2.
    We denote $\textbf{d}(f,P)=\inf\{r\geq 0:d(f,P)\leq r\}$.
It is obvious that $f\in\overline{P}$ if and only if $\textbf{d}(f,P)=0$.
6.1 Tightness
Lemma 6.2.
Let $X$ be an open Whyburn topological space and $(Y,d)$ be metric space, $f\in C(X,Y)$ and suppose that $X$ is $\mathcal{K}_{\Omega}$-Lindelöf. Then for every $P\subseteq Q(X,Y)$ such that $\textbf{d}(f,P)\leq\varepsilon$ there is a subset $S\in[P]^{\leq\omega}$ such that $\textbf{d}(f,S)\leq\varepsilon$.
Proof.
For every $h\in P$ and $n\in\mathbb{N}$ denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in [6]).
Since $\textbf{d}(f,P)\leq\varepsilon$, $\mathcal{V}_{n}=\{V_{h,n}:h\in P\}$ is a semi-open $\omega$-cover of $X$ for every $n\in\mathbb{N}$.
Since $X$ is an open Whyburn topological space, for every $V_{h,n}$ and a finite subset $K$ of $V_{h,n}$ we can consider an open set $W_{K,h,n}$ such that $K\subset\overline{W_{K,h,n}}\subset V_{h,n}$. Then the family $\{W_{K,h,n}:K\in[V_{h,n}]^{<\omega}$ and $h\in P\}\in\mathcal{K}_{\Omega}$ for every $n\in\mathbb{N}$. For each $n\in\mathbb{N}$ we apply that $X$ is $\mathcal{K}_{\Omega}$-Lindelöf and we get the countable family $\{W_{K_{i}(n),h_{i}(n),n}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$. It remains to note that $S=\{h_{i}(n):i,n\in\mathbb{N}\}$ such that $S\in[P]^{\leq\omega}$ and $\textbf{d}(f,S)\leq\varepsilon$.
Lemma 6.3.
Let $X$ be a metrizable $\mathcal{K}_{\Omega}$-Lindelöf space and $F$ be a closed subset of $X$. Then $F$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Proof.
(1) Assume that $IntF=\emptyset$. Since $X$ is Lusin, $F$ is countable and, hence, $F$ is $\mathcal{K}_{\Omega}$-Lindelöf.
(2) Let $IntF\neq\emptyset$. By Theorem 3.6 ($(2)\Rightarrow(5)(a)$) in [5] (Any open subset of $\mathcal{K}_{\Omega}$-Lindelöf space is $\mathcal{K}_{\Omega}$-Lindelöf), $IntF$ is $\mathcal{K}_{\Omega}$-Lindelöf and, by Theorem 3.6 ($(2)\Rightarrow(5)(b)$) (The union of $\mathcal{K}_{\Omega}$-Lindelöf space with countable space is $\mathcal{K}_{\Omega}$-Lindelöf), $F=IntF\cup(F\setminus IntF)$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Theorem 6.4.
For every metrizable space $X$ the following conditions are equivalent:
  1. 1.
    $t(Q_{p}(X,Y))=\omega$ for all metrizable spaces $Y$;
  2. 2.
    There is a metrizable space $Y$ and a function $f\in Q_{p}(X,Y)$ such that $t(f,Q_{p}(X,Y))=\omega$;
  3. 3.
    There is a function $f\in Q_{p}(X,\mathbb{R})$ such that $t(f,Q_{p}(X,\mathbb{R}))=\omega$;
  4. 4.
    $X$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Proof.
$(1)\Rightarrow(2)$ and $(1)\Rightarrow(3)$ are trivial.
$(2)\Rightarrow(4)$. (i) Suppose that $t(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)\neq Y$ where $Y$ is metrizable and $|Y|>1$. Let a family $\gamma=\{V_{\alpha}\}$ of open subsets of $X$ such that $\gamma\in\mathcal{K}_{\Omega}$. Let $y\in Y\setminus f(X)$.
For every $K\in[X]^{<\omega}$ there is $V_{\alpha}\in\gamma$ such that $K=\{x_{1},...,x_{k}\}\subset\overline{V_{\alpha}}$. Since $X$ is metrizable (hence, it is open Whyburn), there are open subsets $W_{1}$,…,$W_{k}$ of $X$ such that $x_{i}\in\overline{W_{i}}\subset V_{\alpha}\cup\{x_{i}\}$ for $i=1,...,k$, $\overline{W_{i}}\cap\overline{W_{j}}=\emptyset$ for $i\neq j$. Consider the function
$f_{K,\alpha}:=\left\{\begin{array}[]{lcl}f(x_{i})\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\overline{W_{i}};\\ y\,\,\,\,\,\,\,\,on\,\,\,X\setminus(\bigcup\{\overline{W_{i}}:i=1,...,k\}).\\ \end{array}\right.$
Let $F=\{f_{K,\alpha}:K\in[X]^{<\omega}\}$. It is clear that $f\in\overline{F}$. Then there is a countably subset $S=\{f_{K_{i},\alpha_{i}}:i\in\mathbb{N}\}$ of $F$ such that $f\in\overline{S}$.
Claim that $\{V_{\alpha_{i}}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$. Let $K\in[X]^{<\omega}$. Consider the open neighborhood $[K,Y\setminus\{y\}]:=\{h\in Q_{p}(X,Y):h(x)\in Y\setminus\{y\}$ for every $x\in K\}$ of the point $f$ in the space $Q_{p}(X,Y)$. Then, there is $f_{K_{i},\alpha_{i}}\in S\cap[K,Y\setminus\{y\}]$. It follows that $K\subset\overline{V_{\alpha_{i}}}$ and, hence, $\{V_{\alpha_{i}}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
(ii) Suppose that $t(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)=Y$ where $Y$ is a metrizable space with a metric $d$.
(1) Claim that there are non-empty open disjoint subsets $U$ and $V$ of $Y$ such that $\overline{Intf^{-1}(U)\cup Intf^{-1}(V)}=X$.
Choose $p,q\in Y$ such that $p\neq q$ and for each $\varepsilon\in(0,d(p,q))$ we consider $U_{\varepsilon}=O_{\varepsilon}(p)$ and $V_{\varepsilon}=X\setminus\overline{O_{\varepsilon}(p)}$. It is clear that $U_{\varepsilon}\cap V_{\varepsilon}=\emptyset$. Note that if $\overline{Intf^{-1}(U_{\varepsilon})\cup Intf^{-1}(V_{\varepsilon})}\neq X$ then $Int(f^{-1}[FrO_{\varepsilon}(p)])\neq\emptyset$. Let us show that for all $\varepsilon$ such a situation cannot exist.
Let’s assume the opposite. For each $\varepsilon\in(0,d(p,q))$ we choose a point $x_{\varepsilon}\in Int(f^{-1}[FrO_{\varepsilon}(p)])$ and a neighborhood $W_{\varepsilon}$ of $x_{\varepsilon}$ such that $\overline{W_{\varepsilon}}\subset Int(f^{-1}[FrO_{\varepsilon}(p)])$. For each $E\subset(0,d(p,q))$ we define the function $h_{E}$:
$h_{E}:=\left\{\begin{array}[]{lcl}p\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\bigcup\{\overline{W_{\varepsilon}}:\varepsilon\in E\};\\ f\,\,\,\,\,\,\,\,on\,\,\,X\setminus(\bigcup\{\overline{W_{\varepsilon}}:\varepsilon\in E\}).\\ \end{array}\right.$
Let $Z=\{h_{E}:E\subset(0,d(p,q))\}$. Note that $Z\subset Q(X,Y)$ and $f=h_{\emptyset}\in Z$. It is clear that $Z$ is homeomorphic to $2^{\mathfrak{c}}$ and, hence, $t(f,Q_{p}(X,Y))>\omega$.
(2) Let $R=Y\setminus(U\cup V)$ where $U$ and $V$ from (1). For each $x\in f^{-1}(R)$ we choose $W_{x}\in\{U,V\}$ such that $f\upharpoonright(f^{-1}[W_{x}]\cup\{x\})$ is quasicontinuous. Let $A=f^{-1}(U)\cup\{x\in f^{-1}(R):W_{x}=U\}$ and $B=f^{-1}(V)\cup\{x\in f^{-1}(R):W_{x}=V\}$. Thus, $X=A\sqcup B$ and $f\upharpoonright A$ and $f\upharpoonright B$ are quasicontinuous.
Let $p\in U$ and $q\in V$.
(3) Let a family $\gamma=\{W_{\alpha}\}$ of open subsets of $X$ such that $\gamma\in\mathcal{K}_{\Omega}$.
We can represent each set $W_{\alpha}$ as a finite disjoint union $W^{1}_{\alpha}\sqcup...\sqcup W^{n}_{\alpha}$ of open sets $W^{j}_{\alpha}$ where $W^{j}_{\alpha}\subset Intf^{-1}(U)$ or $W^{j}_{\alpha}\subset Intf^{-1}(V)$ for each $j=1,...,n$ and $|\overline{W^{j}_{\alpha}}\cap FrA|\leq 1$. Note that since $X$ is an open Whyburn space, we can assume this property.
For each $W_{\alpha}=W^{1}_{\alpha}\sqcup...\sqcup W^{n}_{\alpha}\in\gamma$ and a finite subset $K=\{x_{1},...,x_{n}\}$ of $X$ where $x_{j}\in\overline{W^{j}_{\alpha}}$, $j=1,...,n$, we define the function $h_{W_{\alpha},K}$:
$h_{W_{\alpha},K}:=\left\{\begin{array}[]{lcl}f(x_{j})\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\overline{W^{j}_{\alpha}};\\ q\,\,\,\,\,\,\,\,on\,\,\,A\setminus W_{\alpha}\\ p\,\,\,\,\,\,\,\,on\,\,\,B\setminus W_{\alpha}.\\ \end{array}\right.$
Note that $h_{W_{\alpha},K}\in Q(X,Y)$ and $f\in\overline{\{h_{W_{\alpha},K}:W_{\alpha}\in\gamma\}}$.
Since $t(f,Q_{p}(X,Y))\leq\omega$, there is a countable family $\{h_{W_{\alpha_{i}},K_{i}}:i\in\mathbb{N}\}$ such that $f\in\overline{\{h_{W_{\alpha_{i}},K_{i}}:i\in\mathbb{N}\}}$. It is remain that $\{W_{\alpha_{i}}:i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
$(4)\Rightarrow(1)$. By Lemma 3.1 in [5], $X$ is a Lusin space.
Let $f\in Q_{p}(X,Y)$ and $f\in\overline{P}$ where $P\subset Q_{p}(X,Y)$.
Let us first reduce the general case to the case $Y=\mathbb{H}$. Since $f$ is a quasicontinuous function on a Lusin space, the set $A=f[X]$ is separable. Thus, by Lemma 5.4 there is a function $\varphi:Y\to\mathbb{H}$ such that for every family $T$ of functions $X\to Y$ we have $f\in\overline{T}$ if and only if $f\circ\varphi\in\overline{\{h\circ\varphi:h\in T\}}$, so now we only have to consider the case $Y=\mathbb{H}$.
By Corollary 4.4, there is a $\mathrm{DA}$-family $\mathcal{E}$ for $f$. For every $F\in\mathcal{E}$ denote by $\varepsilon_{F}:=\mathrm{dc}(f,F)$. Thus, for every $F\in\mathcal{E}$ there is $g_{F}\in C(F,Y)$ such that $d(f(x),g_{F}(x))\leq\varepsilon_{F}$ for every $x\in F$.
By Lemma 6.3, $F$ is $\mathcal{K}_{\Omega}$-Lindelöf. Note that $d(g_{F},P\upharpoonright F)\leq\varepsilon_{F}$. By Lemma 6.5, there is $T_{F}\in[P]^{\leq\omega}$ such that $d(g_{F},T\upharpoonright F)\leq\varepsilon_{F}$.
Consider $T=\bigcup\{T_{F}:F\in\mathcal{E}\}$. It is remain note that $T\in[P]^{\leq\omega}$ and $f\in\overline{T}$.
$(3)\Rightarrow(4)$. It is the implication $(2)\Rightarrow(4)$ for $Y=\mathbb{R}$.
6.2 Fan-tightness
Lemma 6.5.
Let $X$ be an open Whyburn topological space and $(Y,d)$ be metric space, $f\in C(X,Y)$ and suppose that $X$ satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$. Then for every $P_{n}\subseteq Q(X,Y)$ such that $\textbf{d}(f,P_{n})\leq\varepsilon$ there is a finite subset $S_{n}$ of $P_{n}$ such that $\textbf{d}(f,\bigcup S_{n})\leq\varepsilon$.
Proof.
For every $n\in\mathbb{N}$ and $h\in P_{n}$ and denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in [6]).
Since $\textbf{d}(f,P)\leq\varepsilon$, $\mathcal{V}_{n}=\{V_{h,n}:h\in P\}$ is a semi-open $\omega$-cover of $X$ for every $n\in\mathbb{N}$.
Since $X$ is an open Whyburn topological space, for every $V_{h,n}$ and a finite subset $K$ of $V_{h,n}$ we can consider an open set $W_{K,h,n}$ such that $K\subset\overline{W_{K,h,n}}\subset V_{h,n}$. Then the family $\{W_{K,h,n}:K\in[V_{h,n}]^{<\omega}$ and $h\in P_{n}\}\in\mathcal{K}_{\Omega}$ for every $n\in\mathbb{N}$.
Since $X$ satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$, for each $n\in\mathbb{N}$ there is a finite family $T_{n}=\{W_{K_{1},h_{1},n},...,W_{K_{s(n)},h_{s(n)}}\}$ such that $\bigcup\{T_{n}:n\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
It remains to note that the family $\{S_{n}:S_{n}=\{h_{1},...,h_{s(n)}\}\subset P_{n},n\in\mathbb{N}\}$ such that $\textbf{d}(f,\bigcup S_{n})\leq\varepsilon$.
Theorem 6.6.
For every metrizable space $X$ the following conditions are equivalent:
  1. 1.
    $vet(Q_{p}(X,Y))=\omega$ for all metric spaces $Y$;
  2. 2.
    There is a metric space $Y$ and a function $f\in Q_{p}(X,Y)$ such that $vet(f,Q_{p}(X,Y))=\omega$;
  3. 3.
    There is a function $f\in Q_{p}(X,\mathbb{R})$ such that $vet(f,Q_{p}(X,\mathbb{R}))=\omega$;
  4. 4.
    $X$ satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$.
Proof.
$(1)\Rightarrow(2)$ and $(1)\Rightarrow(3)$ are trivial.
$(2)\Rightarrow(4)$. (i) Suppose that $vet(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)\neq Y$ where $Y$ is metrizable and $|Y|>1$.
Let $\gamma_{i}=\{V^{i}_{\alpha}\}$ be a family of open subsets of $X$ such that $\gamma_{i}\in\mathcal{K}_{\Omega}$ for each $i\in\mathbb{N}$ and $y\in Y\setminus f(X)$.
For every $i\in\mathbb{N}$ and $K\in[X]^{<\omega}$ there is $V^{i}_{\alpha}\in\gamma_{i}$ such that $K=\{x_{1},...,x_{k}\}\subset\overline{V^{i}_{\alpha}}$. Since $X$ is metrizable (hence, it is open Whyburn), there are open subsets $W_{1}$,…,$W_{k}$ of $X$ such that $x_{s}\in\overline{W_{s}}\subset\overline{V^{i}_{\alpha}}$ for $s=1,...,k$, $\overline{W_{m}}\cap\overline{W_{j}}=\emptyset$ for $m\neq j$. Consider the function
$f^{i}_{K,\alpha}:=\left\{\begin{array}[]{lcl}f(x_{s})\,\,\,\,\,\,\,\,\,\,on\,\,\,\,\overline{W_{s}};\\ y\,\,\,\,\,\,\,\,on\,\,\,X\setminus(\bigcup\{\overline{W_{s}}:s=1,...,k\}).\\ \end{array}\right.$
Let $F_{i}=\{f^{i}_{K,\alpha}:K\in[X]^{<\omega}\}$. It is clear that $f\in\overline{F_{i}}$. Then for every $i\in\mathbb{N}$ there is a finite subset $S_{i}=\{f^{i}_{K_{1},\alpha_{1}},...,f^{i}_{K_{l(i)},\alpha_{l(i)}}\}$ of $F_{i}$ such that $f\in\overline{\bigcup\{S_{i}:i\in\mathbb{N}\}}$.
Claim that $\{V^{i}_{\alpha_{p}}:p\in\{1,...,l(i)\},i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
Let $K\in[X]^{<\omega}$. Consider the open neighborhood $[K,Y\setminus\{y\}]:=\{h\in Q_{p}(X,Y):h(x)\in Y\setminus\{y\}$ for every $x\in K\}$ of the point $f$ in the space $Q_{p}(X,Y)$. Then, there is $i\in\mathbb{N}$ and $p\in\{1,...,l(i)\}$ $f^{i}_{K_{p},\alpha_{p}}\in S\cap[K,Y\setminus\{y\}]$. It follows that $K\subset\overline{V^{i}_{\alpha_{p}}}$ and, hence, $\{V^{i}_{\alpha_{p}}:p\in\{1,...,l(i)\},i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
(ii) If $vet(f,Q_{p}(X,Y))=\omega$ for some $f\in Q_{p}(X,Y)$, $f(X)=Y$ where $Y$ is metrizable and $|Y|>1$. Let $\gamma_{i}=\{W^{i}_{\alpha}\}$ be a family of open subsets of $X$ such that $\gamma_{i}\in\mathcal{K}_{\Omega}$ for each $i\in\mathbb{N}$. Then, similar to implication $(2)\Rightarrow(4)$ in Theorem 6.4, we construct the function $h_{W_{\alpha,i},K}$ for each $W_{\alpha,i}\in\gamma_{i}$.
Let $F_{i}=\{h_{W_{\alpha,i},K}:K\in[X]^{<\omega}\}$. It is clear that $f\in\overline{F_{i}}$. Then for every $i\in\mathbb{N}$ there is a finite subset $S_{i}=\{h^{1}_{W_{\alpha_{1},i},K_{1}},...,h^{n}_{W_{\alpha_{n(i)},i},K_{n(i)}}\}$ of $F_{i}$ such that $f\in\overline{\bigcup\{S_{i}:i\in\mathbb{N}\}}$. It is remain note that $\{W_{\alpha_{p},i}:p\in\{1,...,n(i)\},i\in\mathbb{N}\}\in\mathcal{K}_{\Omega}$.
$(3)\Rightarrow(1)$. It is clear that any space satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$ is $\mathcal{K}_{\Omega}$-Lindelöf. Then, by Lemma 3.1 in [5], $X$ is a Lusin space.
Let $f\in Q_{p}(X,Y)$ and $f\in\overline{P_{n}}$ where $P_{n}\subset Q_{p}(X,Y)$ for each $n\in\mathbb{N}$.
Similar to the previous theorem, now we only have to consider the case $Y=\mathbb{H}$.
By Corollary 4.4, there is a $\mathrm{DA}$-family $\mathcal{E}$ for $f$. For every $F\in\mathcal{E}$ denote by $\varepsilon_{F}:=\mathrm{dc}(f,F)$. Thus, for every $F\in\mathcal{E}$ there is $g_{F}\in C(F,Y)$ such that $d(f(x),g_{F}(x))\leq\varepsilon_{F}$ for every $x\in F$.
By Lemma 6.3, $F$ is $\mathcal{K}_{\Omega}$-Lindelöf.
Note that $d(g_{F},P_{n}\upharpoonright F)\leq\varepsilon_{F}$. By Lemma 6.5, there is $T_{F}\in[P_{n}]^{<\omega}$ such that $d(g_{F},T_{F}\upharpoonright F)\leq\varepsilon_{F}$.
Consider $T=\bigcup\{T_{F}:F\in\mathcal{E}\}$. It is remain note that $f\in\overline{T}$.
$(3)\Rightarrow(4)$. It is the implication $(2)\Rightarrow(4)$ for $Y=\mathbb{R}$.
6.3 Strong fan-tightness
Similar to the previous theorem, the following result is easy to prove.
Theorem 6.7.
For every metrizable space $X$ the following conditions are equivalent:
  1. 1.
    $vet_{1}(Q_{p}(X,Y))=\omega$ for all metric spaces $Y$;
  2. 2.
    There is a metric space $Y$ and a function $f\in Q_{p}(X,Y)$ such that $vet_{1}(f,Q_{p}(X,Y))=\omega$;
  3. 3.
    There is a function $f\in Q_{p}(X,\mathbb{R})$ such that $vet_{1}(f,Q_{p}(X,\mathbb{R}))=\omega$;
  4. 4.
    $X$ satisfies $S_{1}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$.
Combining the results in [4, 5] and this paper, we get the following diagram in the class of metrizable spaces $X$ and $Y$.
$X$ $Q_{p}(X,\,Y)$
countable Fréchet - Urysohn
$S_{1}(K_{\Omega},\,K_{\Omega})$ $\mbox{vet}_{1}(Q_{p}(X,\,Y))=\omega$
$S_{fin}(K_{\Omega},\,K_{\Omega})$ $\mbox{vet}(Q_{p}(X,\,Y))=\omega$
$K_{\Omega}$-Lindelöf $t(Q_{p}(X,\,Y))=\omega$
Diagram 1.
Remark 6.8.
In ([5], Example 6.1) it was shown that, under Jensen’s axiom, there exists a metrizable Lusin space $X$ such that $vet_{1}(Q_{p}(X,\mathbb{R}))=\omega$. An interesting open question (Questions 2 and 3 in [5]) remains:
Is there a $T_{2}$-space $X$ such that $Q_{p}(X,\mathbb{R})$ has countable tightness but none $Q_{p}(X,\mathbb{R})$ has countable (strong) fan-tightness?